OpenStudy (anonymous):

how do you change a standard form to slope intercept form; algebra PLEASE HELP :)

OpenStudy (anonymous):

slope-intercept form is y=mx + b where m is the slope, and b is the y-intercept. Do you have an example problem that we can work on?

OpenStudy (anonymous):

2x+2y=4

OpenStudy (anonymous):

Okay, so basically, all you have to do is solve for y. 2x+2y=4 2y=4-2x y=(4-2x)/2 y=2-x Then make sure that it is in the form y=mx + b y=-x+2 In this case m=-1, and b=2. Remember m is always the coefficient of x.

OpenStudy (anonymous):

why did the signs change from addition to subtraction

OpenStudy (anonymous):

Well, the signs did not change on the actual numbers/variables themselves. Take a closer look. We can rewrite y=2-x as y= (+2)+(-x). This is the same right? Now, because of the Commutative Property of Addition which says that if two numbers are being added together we can switch them around and they will have the same value. So y=(-x)+(+2) which is the same as y=-x+2 You can see that this is true if we use actual numbers. Let's say y=10-2. So y=8 Switch it around like we did above, and we get y=-2+10. y is still equal to 8! So this is okay to do.

OpenStudy (anonymous):

Just make sure that the sign in front of the number stays with that number, and you will be fine.

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

@jkp269 tell me if i did this right 3x+6y=12 changes to y=0.5+2 @jkp269 @jkp269

OpenStudy (anonymous):

First off, there needs to be an x in the equation, but let me figure it out. 6y=12-3x y=(12-3x)/6 y=2-0.5x y=-0.5x+2 Ok, you were almost there, just lost a negative sign, and may have accidentally forgotten to write the x. Remember to SUBTRACT 3x from both sides to get the 6y by itself (I mess this up myself sometimes) and secondly since there is a negative sign in front of 0.5x, you must keep the negative sign in front of it.

OpenStudy (anonymous):

ok yeah what about changing from slope interscept to standard form

OpenStudy (anonymous):

I found this website http://mathforum.org/library/drmath/view/67180.html it takes you through step by step, and does a much better job of explaining it than I could. (My laptop is about to run out of charge) I wish you good luck on your studies!